Shift report 20000815 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Aug 15 2000 - 08:12:34 EDT

  • Next message: http account: "Shift report 20000815 08:00-16:00"

    ------------------------------------------------------------
    Date: Tue Aug 15 08:12:34 2000, Shift: 00:00-08:00
    Supervisor: RW, with: CE
    Log book#4 pages 92 - 94
    
    Shift Summary:
    ------------------------------------------------------------
    1.Beam was cogged at 23:00 and last until 8:00am. The 
    intensity was 4.3e9 (Blue) and 3.0E9 in the beginning
    and 3.0e( and 2.4E9 at 8:00am. The event rate of the 
    trigger 1 was roughly 10 events/sec in the beginning
    abd 2 events/sec at 8:00am. This indicates that the 
    emittance of the beam decreases much faster than the 
    intensity decreases. It may not be a good idea to use
    the same beam for such a long time, if they can 
    reaccecelarate the beam in 3-4 hours, though the 
    operator may not be pleased.
    
    2. Dana changed the HV of TOFW in the day time from 
    -1500V to -1600V.
    
    Run 2254 (MRS at 90deg. with 4KGauss) has completed 
    with 25953 colisions.
    
    Run 2256 has compleyed with 27448 collisions.
    
    Run 2258 has completed at 8:00am
    
    3. During the run2254, DAQ stopped because fb1-evb 
    Error occurs. To restart the run, click Clear Error and
    Begin run. The previous run is automatically ended by
    the Error status. If a new run doesn't start, try again.
    The Error status are not cleared by Clear Error, but don't
    worry about it. 
    
    BBL HV was found off for about 10 channels. We reset the 
    voltages, but thet tripped off again. We left them with
    no HV.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Aug 15 2000 - 08:16:05 EDT